Find the distance between the points (11, 6, 12) and (12, 7, 17). Round to the nearest tenth.
1
Expert's answer
2012-06-15T07:28:53-0400
Distance d=sqrt((12-11)^2+(7-6)^2+(17-12)^2)=sqrt(1^2+1^2+5^2)=sqrt(27)=3sqrt(3) ~5,2